OpenStudy (anonymous):

Terrence normally uses 1/4 teaspoon of dried basil to make his own salad dressing. He decides to change the recipe and include extra dried basil. He decides that the flavor is perfect when including a total of 5/3 teaspoon. Write an addition equation that could be used to determine the additional amount of basil, b, Terrence added. Solve for b. Describe your steps.

OpenStudy (jamesj):

He's going from 1/4 teaspoon to 5/3 teaspoon. The additional amount is the difference between these two: \[ \frac{5}{3} - \frac{1}{4} \] Now simplify that.

OpenStudy (anonymous):

20 - 3/12 = 17/12